# Control-Aware Trajectory Predictions for Communication-Efficient Drone Swarm Coordination in Cluttered Environments

### v1

@@ III-D KKT-informed training of the trajectory prediction neural network
Although such a classical method is widely used, it has two drawbacks. First, it cannot ensure that the estimated planned trajectories serve as a solution to the embedded QP model of DMPC, e.g., whether they satisfy the constraints or can minimize costs. This may lead to estimation errors that can be detrimental to the control performance. Second, this classical method ignores the structural information of the DMPC model, which can be helpful for the training of the neural network. Since UAVs use the same DMPC model, they know perfectly the dynamics and decision processes of other UAVs. Specifically, if UAV $i$ knows all the input of UAV $j$ to the DMPC model, UAV $i$ can accurately obtain the planned trajectories of UAV $j$ simply by solving the DMPC model with UAV $j$’s input.
@@
We will address these two drawbacks by employing a KKT-informed training approach that brings information about the DMPC to the training process. The structure of the training process is illustrated in Figure 2, which involves two parts: (i) a QP generation neural network that generates the parameters of the embedded QP of the DMPC approach and (ii) a differentiable KKT layer for solving QP and backpropagating the gradients.
Figure 3: Example for QP parameter estimation. UAV 1 wants to estimate the planned trajectories of UAV 2, which solves a QP using information about UAVs 3, 4, and 5. Note that UAV 1 can sense the real-time states of UAV 3 but not UAVs 4 and 5. To reconstruct the QP model solved by UAV 2, UAV 1 already has the information about itself and UAV 3 but needs to estimate the parameters of UAV 4 and UAV 5. Further note that in general cases, UAV 1 may not know the number of UAVs that influence the decision-making of UAV 2 (e.g., there could be UAV 6, 7, etc.).
(i) QP generation neural network. This part aims to enable UAV $i$ to estimate the QP parameters to approximate the DMPC model solved by UAV $j$, in particular the QP parameters associated with UAVs beyond the sensing range of UAV $i$. One example is illustrated in Figure 3. To simplify the estimation problem, our QP generation neural network assumes that around each UAV, there are 4 dummy UAVs imposing a significant impact on its decisions. The number 4 is chosen to be the minimum number that can define a 3-dimensional polygon with these UAVs at vertices. The input to the neural network includes the real-time and historical locations $\{\bm{p}_{j^{\prime}}^{k-h}\}_{h=0}^{H}$ of UAV $j$ and its neighbors and obstacles sensed by UAV $i$, along with the estimated planned trajectory produced by either VAE or the EG-based trajectory predictor. The output of the neural network includes the estimated parameters of the QP solved by UAV $j$, which are characterized by the planned trajectories of 4 dummy UAVs at time step $k-1$. The neural network architecture is defined as an EG layer along with fully connected layers.
@@
(2) Differetiable KKT layer. With the estimated QP parameters of UAV $j$, UAV $i$ can solve the QP of UAV $j$ and obtain a primal solution $\tilde{\bm{U}}_{*,j}^{k-1}$ and a dual solution $\tilde{\bm{\lambda}}_{*,j}^{k-1}$. We will then use these solutions to inform the training of the network network parameters. Specifically, the loss function can be written as
$\mathcal{L}_{\text{KKT}}(\tilde{\bm{U}}_{j}^{k-1},\bm{U}_{j}^{k-1})$ $=\alpha||\tilde{\bm{U}}_{j}^{k-1}-\bm{U}_{j}^{k-1}||_{2}^{2}$
$+\beta||\tilde{\bm{U}}_{*,j}^{k-1}-\bm{U}_{j}^{k-1}||_{2}^{2}$ (18)
@@
where the first term aims to minimize the mean square error between the true (i.e., $\bm{U}_{j}^{k-1}$) and estimated values of the planned trajectories produced by VAE and/or the EG-based trajectory predictor (i.e., $\tilde{\bm{U}}_{j}^{k-1}$), and the second term aims to minimize the mean square error between the true trajectory (i.e., $\bm{U}_{j}^{k-1}$) and QP solution ($\tilde{\bm{U}}_{*,j}^{k-1}$).
@@
$\bm{Q}\frac{\partial\bm{w}^{*}}{\partial\bm{u}^{*}}+\frac{\partial\bm{Q}}{\partial\bm{u}^{*}}\bm{w}^{*}+\frac{\partial\bm{q}}{\partial\bm{u}^{*}}+\bm{R}\frac{\partial\bm{\lambda}^{*}}{\partial\bm{u}^{*}}+\frac{\partial\bm{R}}{\partial\bm{u}^{*}}\bm{\lambda}^{*}=\bm{0}$ (21)
$\left(\frac{\partial\bm{\lambda}^{*}}{\partial\bm{u}^{*}}\right)^{T}(\bm{R}\bm{w}^{*}-\bm{b})$
$+\left(\bm{\lambda}^{*}\right)^{T}\left(\frac{\partial\bm{R}}{\partial\bm{u}^{*}}\bm{w}^{*}+\bm{R}\frac{\partial\bm{w}^{*}}{\partial\bm{u}^{*}}-\frac{\partial\bm{b}^{*}}{\partial\bm{u}^{*}}\right)=0$ (22)
Notice that the KKT layer is only used to facilitate the training of the trajectory prediction network instead of as another estimator. This is because the forward propagation of the KKT layer requires solving a sequence of QPs, which can be computationally expensive for real-time implementation. Nevertheless, minimizing the loss function $\mathcal{L}_{\text{KKT}}(\cdot)$ can incorporate the structural information of the QP model into the trained neural network. Hence, this treatment is expected to improve the prediction accuracy.
@@ IV-A Benchmarks
We evaluate our proposed trajectory prediction algorithm by comparing it to the following benchmarks. These seven benchmark algorithms are chosen to evaluate the value of each component of the proposed algorithm.
  • Baseline. The baseline algorithm follows [8]. The input to the algorithm is the real-time and historical location information of all UAVs in the swarm, as well as geometric information about the static obstacles. Notice that this algorithm does not consider the dynamic nature of the UAV alignment. The architecture of the baseline algorithm includes an LSTM-based encoder and an LSTM-based decoder.
@@
  • EG only. This benchmark only involves the EvolveGCN-based trajectory prediction component that uses real-time and historical location information about the surrounding UAVs to estimate their planned trajectories. This benchmark corresponds to the scenario where communication is not available. This benchmark is trained by minimizing the $l_{2}$ loss. The comparison between EG and Baseline shows the value of considering the topological information about the UAV swarm.
@@
  • VAE + KKT. This benchmark evaluates the estimation performance of the VAE component with the aid of the KKT-informed component, where the KKT-informed component is only activated during the training process due to its forward computational complexity.
  • VAE + EG + KKT. The only difference between this benchmark and VAE + EG (Bayesian) is that this benchmark is trained with the differential KKT layer. This benchmark is further evaluated under scenarios with various compression levels, communication frequencies, and packet loss rates. The comparison between VAE + EG + KKT and VAE+EG shows the value of the differentiable KKT layer in the training of the model.
@@ IV-B Implementation Details
[…]
The dynamics of the UAV swarm follow (1) with the control input determined using the DMPC module, where the system matrices of the UAVs (i.e., $\bm{A}_{i}$ and $\bm{B}_{i}$) follow the dynamics of the Crazyflie 2.1 quadrotor with an underlying position controller. The sampling time and DMPC planning horizon are $\Delta T=0.2$ s and $P=16$, respectively. The UAV trajectories within each planning horizon are parameterized as $l=3$ Bezier curves with order $d=5$. For each scenario, we perform a simulation for $T=30$ s. Following [28], we consider Gaussian position measurement noises with zero mean and a default standard deviation of 0.004 m, which will be increased in the testing phase to evaluate the algorithm’s robustness.
We generate the training dataset by randomly generating 300 simulation scenarios according to the aforementioned distributions. Within each scenario, we collect the following information: (1) real-time locations and velocities of each UAV, (2) the location of obstacles, and (3) the QP matrices and solutions of each UAV at each time step. In total, we collected 450,000 samples, which were then transformed into the training data of the neural networks. Note that the training dataset is generated with the benchmark DMPC, meaning that we assume perfect communication channels such that the input to the QP models is always available and accurate.
The experiments are conducted on a computer with an Intel i7-12700 CPU@2.10 GHz and an NVIDIA RTX 3070 GPU. The designed learning networks are implemented in Python using Pytorch 1.13.1. The DMPC control scheme is implemented on MATLAB R2022b. The input size of the VAE module is 16, consistent with the prediction horizon $P$. The input size for all EG modules and LSTM encoder modules is 20, which is the horizon length for UAV past state observation $H$. All hidden layers in the neural network have 128 neurons and the output layer of different modules are all set to be 16. ReLu is selected to be the activate function and all the neural networks are trained under a learning rate of 0.0001.
@@ IV-C Performance in testing scenarios
In this subsection, we show the results of the testing scenarios. Specifically, we generate another set of 50 scenarios for testing, according to the scenario distribution described in Section IV-B. For each testing scenario, we are interested in analyzing control performance and prediction performance. To evaluate control performance, we measure two key factors for each benchmark algorithm: the total incurred cost throughout the simulation period for efficiency and the distances between agents as well as between agents and obstacles for safety. We will further evaluate our algorithm in various scenarios with different parameters to show the robustness of our algorithms in less ideal cases. The data compression levels of VAE, i.e., the latent dimension, range from $L=6$ to $L=42$. Notice that the original dimension of the planned trajectories is $L=48$. The communication frequencies range from $f_{\text{comm}}=5$ Hz to $f_{\text{comm}}=1.25$ Hz. The packet loss rates are set to range from $P_{\text{loss}}=0$ to $P_{\text{loss}}=1/2$, and the measurement noise levels range from 00 to $0.032$. For prediction performance, we will evaluate the mean square errors (MSE) of the predicted planned trajectories compared to the ground-truth planned trajectories solved by DMPC.
TABLE I: Cost comparison between different algorithms
Algorithm
Safe-Agent Cost
Safe-Obs Cost
Control Effort
Cohesion Cost
Total Cost
Baseline
5444
63793
116711
45018
230966
EG
0
0
158302
187283
345586
VAE
0
0
146930
173067
319997
VAE + EG (Bayesian)
0
0
191507
107096
298603
VAE + KKT
0
0
186482
109678
296160
VAE + EG + KKT
0
0
159758
51027
210786
DMPC
0
0
141084
67298
208382
Figure 4: Distance comparison between different algorithms.
We first compare the performance of the benchmarks listed in Section IV-A with all parameters set to the default values, i.e., $L=24$, $f_{\text{comm}}=5$ Hz, $P_{\text{loss}}=0$, and measurement noise level 0.004 m. The resulting average control costs across all scenarios are summarized in Table I, whereby the Safe-Agent Cost penalizes the violation of the safety constraints between UAVs, the Safe-Agent Obs Cost penalizes the violation of the safety constraints between UAVs and obstacles, the Control Effort cost helps to minimize the energy spent to execute the control action, the Cohesion Cost penalizes UAVs being too far away from each other, and the total cost is the sum of all four costs. Figure 4 illustrates agent-agent and agent-obstacle distances, averaged over all testing scenarios, whereas the temporal evolution of these distances in a typical scenario is illustrated in Figures 511. To better understand the control performance, we further summarize the average trajectory prediction errors (in terms of MSE) across all testing scenarios in Tables IIIV.
@@
Value of using real-time location measurement information. We evaluate the value of integrating real-time location information into the prediction of the planned trajectories by comparing VAE and VAE+EG (Bayesian), which shows the value of adding EG. We can see from Table I that both algorithms can ensure safety and prevent collisions since the Safe-Agent Cost and Safe-Obs Cost are zero for both algorithms. We can see this in more detail in Figure 6, Figure 8 that both VAE and VAE+EG (Bayesian) can successfully keep the agent-agent and agent-obstacle distance within the desired threshold. But VAE is dangerously close in terms of agent-agent distance, making it less desirable than VAE+EG (Bayesian). Moreover, it is clear from Table I that VAE+EG (Bayesian) incurs lower costs than VAE. The same holds for prediction performance, as reflected by Tables IIIV. In summary, we can conclude that by adding the EG component, VAE+EG (Bayesian) outperforms VAE. Therefore, with the aid of the Bayesian approach, we can effectively integrate different estimations to enhance the accuracy of planned trajectories.
Value of communication. We evaluate the benefits of having the communicated message by comparing VAE+EG (Bayesian) and EG, which shows the value of having the reconstructed trajectories from VAE. Our conclusion is similar to the above analysis in that VAE+EG (Bayesian) significantly outperforms EG, as demonstrated in Tables IIV, as well as in Figure 4 and Figures 7-8.
@@
Value of the EvolveGCN component in considering time-varying UAV alignment. We demonstrate the value of incorporating the EvolveGCN component by comparing Baseline and EG. From Table I, Figure 4, Figure 5, and Figure 7, we can see that although Baseline outperforms EG in terms of total costs, it cannot ensure safety. It is shown in Figure 5a) and Figure 5b) that there are collisions both between UAVs and between UAVs and the obstacles, and the minimum distance is rather low. The reason is that Baseline does not produce accurate trajectory prediction outcomes, as can be seen from Tables IIIV that the prediction MSE is significantly higher than other benchmarks. This shows that the baseline may not perform well in dynamic environments where the layout of the UAV swarm constantly changes. This also shows the benefits of the EvolveGCN framework in improving safety.

### v2

@@ III-B KKT-informed training of the trajectory prediction neural network
Although such a classical method is widely used, it has two drawbacks. First, it cannot ensure that the predicted trajectories serve as a solution to the embedded QP model of DMPC, e.g., whether they satisfy the constraints or can minimize costs. This may make the predicted trajectories not physically meaningful, leading to estimation errors that can be detrimental to the control performance. Second, this classical method ignores the useful structural information of the DMPC model. Since UAVs use the same DMPC model, they know perfectly the dynamics and decision processes of other UAVs. Specifically, if UAV $i$ knows all the input of UAV $j$ to the DMPC model, UAV $i$ can accurately obtain the planned trajectories of UAV $j$ simply by solving the DMPC model with UAV $j$’s input.
@@
We will address these two drawbacks by employing a KKT-informed training approach that brings information about the DMPC to the training process. The structure of the training process is illustrated in Figure 3. We first instantiate the embedded QP in the DMPC-based controller for UAV $j$, whereby the input is the predicted trajectories of all neighbors of UAV $j$.
@@
With the instantiated QP of UAV $j$, UAV $i$ can solve the QP of UAV $j$ and obtain a primal solution $\tilde{\bm{U}}_{*,j}^{k}$ and a dual solution $\tilde{\bm{\lambda}}_{*,j}^{k}$. Ideally, the primal solution $\tilde{\bm{U}}_{*,j}^{k}$ should be close to the predicted trajectory $\tilde{\bm{U}}_{j}^{k}$ calculated by the trajectory predictor. We leverage this insight via a KKT-informed training approach, which uses the primal solution $\tilde{\bm{U}}_{*,j}^{k}$ and dual solution $\tilde{\bm{\lambda}}_{*,j}^{k}$ to inform the training of the network parameters. Specifically, we aim to minimize the loss function written as
$\mathcal{L}_{\text{KKT}}(\tilde{\bm{U}}_{j}^{k},\bm{U}_{j}^{k})$ $=\alpha||\tilde{\bm{U}}_{j}^{k}-\bm{U}_{j}^{k}||_{2}^{2}+\beta||\tilde{\bm{U}}_{j}^{k}-\tilde{\bm{U}}_{*,j}^{k}||_{2}^{2}$ (11)
@@
where the first term aims to minimize the mean square error between the true (i.e., $\bm{U}_{j}^{k}$) and estimated values of the planned trajectories produced the EG-based trajectory predictor (i.e., $\tilde{\bm{U}}_{j}^{k}$), and the second term aims to minimize the mean square error between the QP solution (i.e., $\tilde{\bm{U}}_{*,j}^{k}$) and the estimated planned trajectory (i.e., $\tilde{\bm{U}}_{j}^{k}$).
@@
$\bm{Q}\frac{\partial\bm{w}^{*}}{\partial\bm{u}^{*}}+\frac{\partial\bm{Q}}{\partial\bm{u}^{*}}\bm{w}^{*}+\frac{\partial\bm{q}}{\partial\bm{u}^{*}}$
$+\bm{R}\frac{\partial\bm{w}^{*}}{\partial\bm{u}^{*}}+\frac{\partial\bm{R}}{\partial\bm{u}^{*}}\bm{w}^{*}+\bm{G}\frac{\partial{\lambda}^{*}}{\partial\bm{u}^{*}}+\frac{\partial\bm{G}}{\partial\bm{u}^{*}}{\lambda}^{*}={0}$ (13a)
$\frac{\partial\bm{R}}{\partial\bm{u}^{*}}\bm{w}^{*}+\bm{R}\frac{\partial\bm{w}^{*}}{\partial\bm{u}^{*}}-\frac{\partial\bm{b}^{*}}{\partial\bm{u}^{*}}=0$ (13b)
$\left(\frac{\partial{\lambda}^{*}}{\partial\bm{u}^{*}}\right)^{T}(\bm{G}\bm{w}^{*}-\bm{h})$
$+\left({\lambda}^{*}\right)^{T}\left(\frac{\partial\bm{G}}{\partial\bm{u}^{*}}\bm{w}^{*}+\bm{G}\frac{\partial\bm{w}^{*}}{\partial\bm{u}^{*}}-\frac{\partial\bm{h}^{*}}{\partial\bm{u}^{*}}\right)=0$ (13c)
Notice that the KKT layer is used solely to facilitate the training of the trajectory prediction network rather than as an additional estimator. This is because the forward propagation of the KKT layer requires solving a sequence of QPs, which can be computationally expensive for real-time implementation. In this way, the proposed KKT-informed training process can be applied to different trajectory predictor settings (e.g., with or without VAE-based compression and reconstruction module). Nevertheless, minimizing the loss function $\mathcal{L}_{\text{KKT}}(\cdot)$ can incorporate the structural information of the QP model into the trained neural network. Hence, this treatment is expected to improve the accuracy of the prediction.
@@ IV-A Benchmarks
We evaluate our proposed trajectory prediction algorithm by comparing it to the following benchmarks. These nine benchmark algorithms are chosen to evaluate the value of each component of the proposed algorithm.
  • LIA. This benchmark follows [21], which is a learning interaction-aware (LIA) trajectory prediction algorithm dedicatedly designed for UAV swarm trajectory prediction task. The input to the algorithm is the real-time and historical location information of all UAVs in the swarm, as well as geometric information about the static obstacles. Notice that this algorithm does not consider the dynamic nature of the UAV alignment. The architecture of the LIA algorithm includes an LSTM-based encoder and an LSTM-based decoder.
  • Y-Net. This benchmark follows [25], a state-of-the-art trajectory prediction algorithm that handles the uncertainty of future trajectories with factorized goal and path multimodalities, which is widely recognized in various trajectory prediction tasks [40, 41]. We extend this work from its original 2D domain ($x$ and $y$) to our 3D domain ($x$, $y$, and $z$). The input to the algorithm consists of real-time and historical location information of predicted UAVs in the swarm, as well as map information about static obstacles. UAV dynamics are not considered in this benchmark. The architecture of this benchmark includes a Convolutional Neural Network (CNN)-based encoder and two CNN-based decoders.
  • EG only. This benchmark only involves the EvolveGCN-based trajectory prediction component that uses real-time and historical location information about the surrounding UAVs to estimate their planned trajectories. This benchmark corresponds to the scenario where communication is not available. This benchmark is trained by minimizing the $l_{2}$ loss. The comparison between EG and LIA/Y-Net shows the value of considering the topological information about the UAV swarm.
@@
  • EG + KKT. This benchmark evaluates the estimation performance of the EG component with the aid of the KKT-informed component, where the KKT-informed component is only activated during the training process due to its forward computational complexity. The comparison between EG + KKT and EG shows the value of the differentiable KKT layer in the training of the model.
  • VAE + KKT. This benchmark evaluates the estimation performance of the VAE component with the aid of the KKT-informed component, and the KKT-informed component is only activated during the training process.
  • VAE + EG + KKT. The only difference between this benchmark and VAE + EG (Bayesian) is that this benchmark is trained with the differential KKT layer. This benchmark is further evaluated under scenarios with various compression levels, communication frequencies, and packet loss rates. The comparison between VAE + EG + KKT and VAE+EG (Bayesian) shows the value of the differentiable KKT layer in the training of the model.
@@ IV-B Implementation details
The swarm size is sampled uniformly at random from set $N=\{n\in\mathbb{Z}|6\leq n\leq 14\}$. The initial locations of the UAVs are generated within a space $\{(p_{x},p_{y},p_{z})|p_{x}\in[0,3],p_{y}\in[-4,4],p_{z}\in[-1,1]\}$ (in meters) using the Poisson Disk Sampling approach, which ensures that the distance between any pair of UAVs is above a minimum Euclidean distance of 0.7 m. The minimum, maximum, and desired velocities of each UAV are assumed to be -1.5 m/s, 1.5 m/s, and 0.5 m/s, respectively. The minimum and maximum acceleration rates are assumed to be -1 m/s2 and 1 m/s2, respectively. We aim to prevent collisions between UAV $i$ and other UAVs/obstacles by requiring $||\bm{p}_{i}-\bm{p}||_{\bm{E}}\geq r_{\min}$, where $\bm{E}$ represents a linear transformation matrix, $\bm{p}_{i}$ represents the location of UAV $i$, $\bm{p}$ represents the location of another UAV or the closest point of an obstacle, and $r_{\min}$ represents a safety buffer chosen as $r_{\min}=0.15$ m. If $||\bm{p}_{i}(k)-\bm{p}(k)||_{\bm{E}}<r_{\text{coll}}$, we say that collisions occur, where the collision threshold $r_{\text{coll}}$ is chosen as 0.07 m.
The dynamics of the UAV swarm follow Eq.(1) with the control input determined using the DMPC module, where the system matrices of the UAVs (i.e., $\bm{A}_{i}$ and $\bm{B}_{i}$) follow the dynamics of the Crazyflie 2.1 quadrotor with an underlying position controller. The sampling time and DMPC planning horizon are $\Delta T=0.2$ s and $P=16$, respectively. The UAV trajectories within each planning horizon are parameterized as $l=3$ Bézier curves with order $d=5$. The weights of objective function Eq. (2a), $q_{\text{mig}},l_{\text{saf}},q_{\text{saf}},l_{\text{coh}},q_{\text{coh}},q_{\text{eft}}$, are set to be $701,1e4,1e2,1e3,10,$ 70 respectively. For each scenario, we perform a simulation for $T=30$ s. Following [17], we consider Gaussian position measurement noises with zero mean and a default standard deviation of 0.004 m, which will be increased in the testing phase to evaluate the algorithm’s robustness.
We generate the training dataset by randomly generating 300 simulation scenarios according to the aforementioned distributions. Within each scenario, we collect the following information: (1) real-time locations and velocities of each UAV, (2) the location of obstacles, and (3) the QP matrices and solutions of each UAV at each time step. In total, we collected 450,000 samples, which were then transformed into the training data of the neural networks. Note that the training dataset is generated with the benchmark DMPC, meaning that we assume perfect communication channels such that the input to the QP models is always available and accurate. We only collected data from episodes in which the swarm successfully reached the goal, as this implies the successful achievement of the control aim.
The experiments are conducted on a computer with an Intel i7-12700 CPU@2.10 GHz and an NVIDIA RTX 3070 GPU. The designed learning networks are implemented in Python using Pytorch 1.13.1. The DMPC control scheme is implemented on MATLAB R2022b and the UAV swarm is simulated using Swarmlab [42]. The input size of the VAE module is 16, consistent with the prediction horizon $P$. The input size for all EG modules and LSTM encoder modules is 20, which is the horizon length for UAV past state observation $H$. All hidden layers in the neural network have 128 neurons and the output layer of different modules are all set to be 16. ReLu is selected to be the activate function and all the neural networks are trained under a learning rate of 0.0001.
@@ IV-C Performance in testing scenarios
In this subsection, we show the results of the testing scenarios. Specifically, we generate another set of 50 scenarios for testing, according to the scenario distribution described in Section IV-B. For each testing scenario, we are interested in analyzing control performance and prediction performance. To evaluate control performance, we measure two key factors for each benchmark algorithm: the total incurred cost throughout the simulation period for efficiency and the distances between agents as well as between agents and obstacles for safety. We will further evaluate our algorithm in various scenarios with different parameters to show the robustness of our algorithms in less ideal cases. The data compression levels of VAE, i.e., the latent dimension, range from $L=6$ to $L=42$. Notice that the original dimension of the planned trajectories is $L=48$. The communication frequencies range from $f_{\text{comm}}=5$ Hz to $f_{\text{comm}}=1.25$ Hz. The packet loss rates are set to range from $P_{\text{loss}}=0$ to $P_{\text{loss}}=1/2$, and the measurement noise levels range from 00 to $0.032$. For prediction performance, we will evaluate the mean square errors (MSE) of the predicted planned trajectories compared to the ground-truth planned trajectories solved by DMPC. During the testing phase, the UAV swarm consistently reaches its goal under different benchmarks.
TABLE I: Cost comparison between different algorithms, with costs in units of $10^{5}$
Algorithm Safe-Agent Cost Safe-Obs Cost Migra-tion Cost Control Effort Cohesion Cost Total Cost
LIA 0.0544 0.638 1.09 1.17 0.450 3.40
Y-Net 0 0 1.85 2.04 2.38 6.27
EG 0 0 1.19 1.58 1.87 4.65
VAE 0 0 1.17 1.42 1.79 4.39
EG + KKT 0 0 1.75 1.91 0.544 4.20
VAE + EG (Bayesian) 0 0 1.22 1.86 1.10 4.18
VAE + KKT 0 0 1.15 1.34 0.971 3.46
VAE + EG + KKT 0 0 1.18 1.51 0.615 3.30
DMPC 0 0 1.16 1.41 0.673 3.24
((a)) Agent-agent distance comparison of different benchmarks.
((b)) Minimum agent-obstacle distance of different benchmarks.
Figure 4: Safe distance comparison of different benchmarks.
We first compare the performance of the benchmarks listed in Section IV-A with all parameters set to the default values, i.e., $L=24$, $f_{\text{comm}}=5$ Hz, $P_{\text{loss}}=0$, and measurement noise level 0.004 m. The resulting average control costs across all scenarios are summarized in Table I, whereby the Safe-Agent Cost penalizes the violation of the safety constraints between UAVs, the Safe-Agent Obs Cost penalizes the violation of the safety constraints between UAVs and obstacles, the Migration Cost penalizes the distance to the pre-defined migration point, the Control Effort cost helps to minimize the energy spent to execute the control action, the Cohesion Cost penalizes UAVs being too far away from each other, and the total cost is the sum of all five costs. Figure 4 illustrate agent-agent and agent-obstacle distances, averaged over all testing scenarios, whereas the temporal evolution of these distances in a typical scenario is illustrated in Figure 5. Table II summarizes the average trajectory prediction errors (measured as MSE along the $x$, $y$, and $z$ axes using the Euclidean norm) across all testing scenarios.
@@
Value of the EvolveGCN component in considering time-varying UAV alignment. We demonstrate the value of incorporating the EvolveGCN component by comparing LIA, Y-Net and EG. From Table I, Figure 4, Figure 5a) – Figure 5f), we can see that although LIA outperforms Y-Net and EG in terms of total costs, it cannot ensure safety. It is shown in Figure 5a) and  5b) that there are collisions both between UAVs and between UAVs and the obstacles. Moreover, we find that Y-Net maintains the largest cost in the migration cost, control effort and cohesion cost, consequently resulting in the largest total cost. It is shown in Figure 5c) and 5d) that the maximum distance between agents is high, while the minimum distance between agents and obstacles is rather low. The reason can be attributed to the fact that LIA and Y-Net do not produce accurate trajectory prediction outcomes, as can be seen in Table II where the prediction MSE is higher than other benchmarks. This indicates that LIA and Y-Net may not perform well in dynamic environments where the layout of the UAV swarm constantly changes. This also highlights the benefits of the EvolveGCN framework in improving prediction accuracy.